# From Freshness to Effectiveness: Goal-Oriented Sampling for Remote Decision Making

### v2

@@ 24            $\lambda_{\downarrow}=\lambda$;
Algorithm 2 A Three-layer Algorithm (A Variant of [85, Section IV.C])
@@ 6      Solve LP Problem 7 to obtain $Q^{\star}(f_{\max})$;
In the above three-layer algorithm, each update of $\lambda$ or $\theta$ necessitates solving the inner-layer MDP using the $\tau$-RVI algorithm. This process incurs a high computational complexity due to the repeated execution of the $\tau$-RVI algorithm required to iteratively search and optimize the parameters $\lambda$ and $\theta$.
@@ VII-A QuickBLP
($i$). If the root of $U(\rho)$, denoted as $\rho^{\star}$, satisfies $\mathcal{F}^{(\rho^{\star})^{-}}\geq 1/f_{\max}$, then the optimal value of Problem 1 is $h^{\star}=\rho^{\star}$, and the optimal policy for Problem 1 is $\phi_{\rho^{\star}}^{\star}$, as defined in (IV-C1);
($ii$). If $\mathcal{F}^{(\rho^{\star})^{-}}<1/f_{\max}$, the optimal value of Problem 1 is
@@
Theorem 8 leads to a one-layer two-stage algorithm presented in Algorithm 3. In the first stage, the algorithm solves the unconstrained MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ and determines the root of $U(\lambda)$, denoted by $\rho^{\star}$. This root is obtained by implementing OnePDSI in Iteration 3. If the condition specified in part ($i$) of Theorem 8 holds true, then the root $h^{\star}$ is immediately found as $h^{\star}=\rho^{\star}$. If the condition of part ($i$) of Theorem 8 is not met, the algorithm proceeds to the second stage. Here, the LP problem [^11] Once an LP problem is established, we can say the solution is achieved. This is based on the mature techniques such as the simplex algorithm and the interior point method for solving LP problems. Problem 7 is solved. The solution of this LP provides the optimal value $Q^{\star}(f_{\max})$. Finally, the root $h^{\star}$ is explicitly determined as $h^{\star}=f_{\max}\cdot Q^{\star}(f_{\max})$. Compared to the previous Algorithm 2, this newly proposed algorithm eliminates the need for multiple executions of the $\tau$-RVI for searching $h^{\star}$ and $\theta^{\star}_{h^{\star}}$. Table II compares the two algorithms.
VII-B Sensitivity Analysis and Sampling Frequency Threshold
@@
To better understand the relationship between $h^{\star}$ and the maximum rate constraint parameter $f_{\max}$, we conduct a sensitivity analysis in this subsection to explore how $h^{\star}$ depends on the maximum rate constraint with parameter $f_{\max}$. The key result is stated in the following Lemma, whose proof can be found in Appendix LABEL:apposa.
@@
(Sensitivity Analysis). Define $f_{\max}^{T}$ as:
$f_{\max}^{T}\triangleq\frac{1}{\mathcal{F}^{(\rho^{\star})^{-}}},$ (59)
the following assertions hold true:
($i$). If $f_{\max}\geq f_{\max}^{T}$, $h^{\star}$ is independent of $f_{\max}$;
($ii$). If $f_{\max}<f_{\max}^{T}$, $h^{\star}=f_{\max}Q^{\star}(f_{\max})$ is monotonically non-increasing with $f_{\max}$, with the derivative given as:
$\frac{\mathrm{d}h^{\star}}{\mathrm{d}f_{\max}}=U(\lambda^{\star}),$ (60)
where
[…]
@@
TABLE II: Comparing Algorithm 2 and Algorithm 3
Aspect Algorithm 2 [85, Section IV.C] Algorithm 3 (Proposed QuickBLP)
Loop Structure Three nested layers Single-layer with two stages
Number of Iterations $\log\frac{1}{\epsilon_{1}}\cdot\log\frac{1}{\epsilon_{2}}$ $\tau$-RVI calls 1 OnePDSI call + 1 optional LP call
(LP is triggered only if $\mathcal{F}^{(\rho^{\star})^{-}}<1/f_{\max}$)
Parameter Sensitivity Depends on tolerance parameter $\epsilon_{1}$, $\epsilon_{2}$ Parameter-free (Solves $h^{\star}$ directly)
Note: Both $\tau$-RVI and OnePDSI have exponential convergence rates where the optimality gap decreases exponentially in the worst case.
@@ VIII-A1 Sampling Policies
In this paper, the sampling policy that is co-designed with the goal-oriented remote decision making to minimize the long-term average cost in Problem 1 is referred to as “Goal-oriented sampling”. We compare this sampling policy with the following benchmarks:
  • Uniform sampling: The sampling is activated periodically with a constant sampling interval $d\in\mathbb{N}^{+}$. In this case, the sampling process follows $S_{i+1}-S_{i}=d$ for $\forall i$ (see [4, Section VI] for a detailed system description of uniform sampling with queuing and random service delay). Since the sampling interval $d$ is limited to integer values in the discrete-time MDP setting, the corresponding sampling frequency $1/d$ also takes discrete values. Hence, the simulation yields a performance curve comprising discrete points, as shown in Fig. 8.
@@
  • AoI-optimal sampling: The AoI-optimal policy determines $Z_{i}$ by [63, Theorem 4], which is a threshold-based policy $Z_{i}=\max(0,\beta-Y_{i})$, where $\beta$ is the solution to the following equations:
@@
    with
    $z(Y)=\max(0,\beta-Y).$ (63)
    [63, Algorithm 2] presents a low-complexity bisection search algorithm to solve $\beta$.
@@ VIII-A2 Remote Decision Making
We consider the following decision making policies:
@@
  • Long-term optimal decision policy $\pi^{\star}$: The long-term optimal policy thinks ahead to minimize the long-term average cost over time. We compute it by solving the ACOE of the primal MDP using the RVI algorithm. The resulting policy employed at the remote decision maker $\pi^{\star}:\mathcal{S}\rightarrow\mathcal{A}$ selects an action based on the latest available observation $X_{t-\Delta(t)}$, according to $a_{t}=\pi^{\star}(X_{t-\Delta(t)})$. In our experiments and analysis, we use this decision-making policy as the default for the benchmark sampling policies.
Figure 7: Average cost vs. Average Delay $\mathbb{E}[Y_{i}]$.
Figure 8: Average cost vs. Maximum sampling frequency $f_{\max}$.
@@ VIII-B1 Primal MDP
Figure 9: Simulation Setup: Transition diagram of the primal MDP given in Appendix LABEL:appendixh.
As a case study, we specifically consider a benchmark setup for clarity and insight. In this setup, the sizes of the state space and the action space of the primal MDP are both $2$, consisting of states $s_{0},s_{1}$ and actions $a_{0},a_{1}$ [^12] The binary state space is broadly representative of many real-world systems. For instance, it can model the occurrence or absence of critical events such as fires, industrial failures, security breaches, or anomalies in equipment status. The binary action space can be interpreted as whether or not to apply a control intervention in response to the critical event, e.g. extinguishing a detected fire or repairing a malfunctioning piece of equipment.. The primal MDP tuple is detailed in Appendix LABEL:appendixh, and is visualized by transition diagram given in Fig. 9.
@@
indicating a constant action regardless of state. Consequently, this decision making policy does not utilize state information at the remote decision maker, and thus the value of information transmission is null in this context.
@@
which clearly depends on the system state. In this case, the remote decision maker must rely on the potentially stale state to select the “right” action, highlighting the utility of information transmission under the decision making policy.
VIII-B2 Random Delay Distribution
@@
As a representative example, we consider an i.i.d. Bernoulli random delay model in which each delay $Y_{i}$ takes value $1$ with probability $p$ and $11$ with probability $1-p$, i.e., $\Pr(Y_{i}=1)=p,\Pr(Y_{i}=11)=1-p,\forall i\in\mathbb{Z}^{+}$. This binary-delay formulation enables tractable analysis while capturing key aspects of stochastic delay. The framework, however, is not limited to this case and can be generalized to arbitrary delay distributions.
VIII-C Discussions and Insights
Fig. 7 illustrates the performance comparison between various sampling and decision-making policies under different average delays $\mathbb{E}[Y_{i}]$. The proposed goal-oriented sampling consistently outperforms all benchmark methods across the full range of average delay values. A key observation is that the myopic decision-making policy, which always selects a fixed action regardless of state information, yields a constant cost independent of delay. This highlights an insight: in scenarios where the control policy is insensitive to state information, enhancing the communication channel (e.g., by reducing delay) does not lead to any improved goal-oriented decision making performance.

### v3

@@ 24    $\lambda_{\downarrow}=\lambda$;
Algorithm 2 A Three-layer Algorithm (A Variant of [110, Section IV.C])
@@ 6  Solve LP Problem 7 to obtain $Q^{\star}(f_{\max})$;
In the three-layer algorithm, each update of $\lambda$ or $\theta$ necessitates solving the inner-layer MDP using the $\tau$-RVI algorithm. This process incurs a high computational complexity due to the repeated execution of the $\tau$-RVI algorithm required to iteratively search and optimize the parameters $\lambda$ and $\theta$.
@@ VII-A QuickBLP
(ii). If the root of $U(\rho)$, denoted as $\rho^{\star}$, satisfies $\mathcal{F}^{(\rho^{\star})^{-}}\geq 1/f_{\max}$, then the optimal value of Problem 1 is $h^{\star}=\rho^{\star}$, and the optimal policy for Problem 1 is $\phi_{\rho^{\star}}^{\star}$, as defined in (IV-B1);
(iiii). If $\mathcal{F}^{(\rho^{\star})^{-}}<1/f_{\max}$, the optimal value of Problem 1 is
@@
Theorem 8 leads to a one-layer two-stage algorithm presented in Algorithm 3. In the first stage, the algorithm solves the unconstrained MDP $\mathscr{P}_{\text{MDP}}(\lambda)$ and determines the root of $U(\lambda)$, denoted by $\rho^{\star}$. This root is obtained by implementing OnePDSI in Iteration 3. If the condition specified in part ($i$) of Theorem 8 holds true, then the root $h^{\star}$ is immediately found as $h^{\star}=\rho^{\star}$. If the condition of part ($i$) of Theorem 8 is not met, the algorithm proceeds to the second stage. Here, the LP problem [^12] Once an LP problem is established, it can be solved to global optimality using well-established algorithms such as the simplex and modern primal–dual interior-point methods. For problems whose variable set is large, mature decomposition techniques such as column generation allow solving the LP without ever forming the entire matrix, while preserving exact optimality. Problem 7 is solved. The solution of this LP provides the optimal value $Q^{\star}(f_{\max})$. Finally, the root $h^{\star}$ is explicitly determined as $h^{\star}=f_{\max}\cdot Q^{\star}(f_{\max})$. Compared to the previous Algorithm 2, this newly proposed algorithm eliminates the need for multiple executions of the $\tau$-RVI for searching $h^{\star}$ and $\theta^{\star}_{h^{\star}}$.
VII-B Polynomial Complexity Results
In this subsection, we analyze that both the methods proposed in this paper are polynomial in time and space complexity. For the three-layer Algorithm 2, the computational cost of a single $\tau$-RVI iteration is $\mathcal{O}(|\mathcal{Z}||\mathcal{S}|^{2}|\mathcal{Y}|^{2}|\mathcal{A}|^{3})$, and according to Theorem 3, the $\tau$-RVI requires at most $\mathcal{O}\left(\log(1/\delta)\right)$ iterations to reach accuracy $\delta$. Consequently, the time complexity of the three-layer Algorithm 2 is
$\mathcal{O}\left(\log(\frac{1}{\epsilon_{1}})\log(\frac{1}{\epsilon_{2}})\log(\frac{1}{\delta})|\mathcal{Z}||\mathcal{S}|^{2}|\mathcal{Y}|^{2}|\mathcal{A}|^{3}\right),$ (63)
which is polynomial in all problem parameters. In terms of space complexity, the algorithm stores one value vector $\tilde{V}_{K}(\gamma;\lambda)$ of size $\mathcal{O}(|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|)$ together with the sparse transition structure of size $\mathcal{O}(|\mathcal{Z}||\mathcal{S}|^{2}|\mathcal{Y}|^{2}|\mathcal{A}|^{3})$. However, the overall computational cost can be dominated by the outer layers: the three-layer algorithm performs a large number of nested loops to reach accuracies $\epsilon_{1},\epsilon_{2}$ and $\delta$. When these tolerances are set small, the number of outer loops can be very high, and the resulting run time, though still polynomial, might become impractical.
In contrast, the proposed QuickBLP addresses this challenge by eliminating the outer-loop search. In the first stage, OnePDSI incurs $\mathcal{O}(|\mathcal{Z}||\mathcal{S}|^{2}|\mathcal{Y}|^{2}|\mathcal{A}|^{3})$ per iteration and converges in $\mathcal{O}\!\big(\log(1/\delta)\big)$ iterations (Theorem 6), i.e., the complexity of OnePDSI is polynomial:
$\mathcal{O}\left(\log(\frac{1}{\delta})|\mathcal{Z}||\mathcal{S}|^{2}|\mathcal{Y}|^{2}|\mathcal{A}|^{3}\right).$ (64)
If the stopping condition of Theorem 8 is satisfied, the algorithm terminates, and no LP needs to be solved. Only when this condition is not met does the second stage become active, in which the LP Problem 7 with $n=|\mathcal{Z}|\times|\mathcal{S}|\times|\mathcal{Y}|\times|\mathcal{A}|^{2}$ variables and $m=|\mathcal{S}\times\mathcal{Y}\times\mathcal{A}|+2$ equality constraints must be solved. The LP is handled via the Column Generation (CG) method [26] to decompose the Master Problem (MP) into some manageable Restricted Master Problems (RMP). Each RMP is solved using a primal–dual interior-point method [15, Chap. 11]. Since the occupancy-measure LP admits optimal basic feasible solutions, the solution vector $\mathbf{x}$ can be chosen with at most $m$ nonzeros; hence, at CG iteration $t$ the RMP involves $n_{t}\leq m$ active columns. A primal–dual interior-point method requires $\mathcal{O}(\sqrt{m}\log(1/\epsilon))$ iterations [33, Theo. 3.1], each dominated by solving a sparse Schur system with $n_{t}$ variables and $m$ constraints, whose factorization cost $T_{\text{fact}}(m,n_{t})$ ranges from $\mathcal{O}(m^{1.5})$ to $\mathcal{O}(m^{2})$. As a result, the total cost for solving the LP via CG and primal–dual interior-point methods can be upper bounded by:
$\mathcal{O}\left(\sum_{t=1}^{q}\sqrt{m}\log(1/\epsilon)\cdot T_{\text{fact}}(m,n_{t})\right),$ (65)
where $q\leq m$ is the number of CG iterations and $\sum_{t=1}^{q}\leq m^{2}$. This yields a worst-case complexity between $\mathcal{O}(m^{3.5})$ to $\mathcal{O}(m^{4})$. The space complexity to solve the LP is dominated by storing and factorizing the Schur complement system of order $m$. Since each RMP contains $n_{t}\leq m$ active columns, the overall memory requirement is $\mathcal{O}(m)$ to $\mathcal{O}(m^{2})$, depending on the sparsity structure of the constraints.
Discussion: Both the algorithms achieve polynomial complexity. The three-layer algorithm admits a lower-degree polynomial bound in theory but suffers from heavy nested searches, which can make the runtime large in practice. In contrast, QuickBLP adopts a two-stage design: in many instances the first stage alone suffices, and the second-stage LP is only occasionally invoked. Although the LP stage has a higher worst-case bound, the combination of CG and sparse interior-point methods makes the practical complexity often much lower than the theoretical worst case. As a result, QuickBLP is typically far more efficient in real problem instances.
VII-C Approximate LP: Scaling to Large Spaces
In large-scale systems, the growth of the state space $\mathcal{S}$, the action space $\mathcal{A}$, and the delay space $\mathcal{Y}$ renders optimally solving the problem intractable. A key advantage of QuickBLP is that its linear programming formulation naturally accommodates approximate linear programming (ALP), making it well suited for scaling to large state–action–delay spaces. Following [60], we first adopt a linear approximation of the occupation measure $x(\gamma,z,a)$ to reduce the number of variables:
$x(\gamma,z,a)$ $\approx\hat{x}(\gamma,z,a;\boldsymbol{\theta})$ (66)
$\triangleq\mu_{0}(\gamma,z,a)+\sum_{i=1}^{d}\theta_{i}\psi_{i}(\gamma,z,a)$
$=\mu_{0}(\gamma,z,a)+\boldsymbol{\theta}\boldsymbol{\psi}^{T},$
where $\mu_{0}(\gamma,z,a)$ is a fixed baseline function, $\{\psi_{i}(\gamma,z,a)\}_{i=1}^{d}$ are predefined basis functions, and $\boldsymbol{\theta}\in\mathbb{R}^{d}$ is the parameter vector to be optimized. This reduces the number of variables from $n=|\mathcal{Z}|\times|\mathcal{S}|\times|\mathcal{Y}|\times|\mathcal{A}|^{2}$ to feature dimension $d$. Substituting this approximation into the LP Problem 7 yields the following ALP:
Problem 8 (Linear Programming Approximation).
$Q^{\star}(f_{\max})\approx\min_{\boldsymbol{\theta}}\sum_{\gamma,z,a}q(\gamma,z,a)(\mu_{0}(\gamma,z,a)+\boldsymbol{\theta}\boldsymbol{\psi}^{T})$ (67a)
$\mathrm{s.t.}\quad$ $\sum_{\gamma,z,a}f(z)\hat{x}(\gamma,z,a;\boldsymbol{\theta})=\frac{1}{f_{\max}},$ (67b)
$\sum_{z,a}\hat{x}(\gamma^{\prime},z,a;\boldsymbol{\theta})=\sum_{\gamma,z,a}p(\gamma^{\prime}|\gamma,z,a)\hat{x}(\gamma,z,a;\boldsymbol{\theta}),$
$\sum_{\gamma,z,a}\hat{x}(\gamma,z,a;\boldsymbol{\theta})=1,$ (67d)
$\hat{x}(\gamma,z,a;\boldsymbol{\theta})\geq 0,~~\forall\gamma\in\mathcal{S}\times\mathcal{Y}\times\mathcal{A},z\in\mathcal{Z},a\in\mathcal{A},$ (67e)
The above ALP reduces the number of decision variables from $n=|\mathcal{Z}|\times|\mathcal{S}|\times|\mathcal{Y}|\times|\mathcal{A}|^{2}$ to a fixed feature dimension $d$, but the number of equality constraints still scales with the sizes of $\mathcal{S}$, $\mathcal{Y}$, and $\mathcal{A}$. This residual dependence on the system size limits the scalability of the approximation unless the constraints are further relaxed.
To achieve scalability, the hard constraints can be relaxed into soft penalties in the objective function [60]. Specifically, the objective function is augmented with weighted $\ell_{1}$ violations of the constraints, yielding a surrogate loss. This reformulation converts the constrained LP into an unconstrained stochastic convex optimization problem over the low-dimensional parameter $\boldsymbol{\theta}\in\mathbb{R}^{d}$, which can be solved efficiently using stochastic subgradient descent. As a result, the overall complexity scales only with the feature dimension $d$, becoming independent of $\mathcal{S}$, $\mathcal{Y}$, and $\mathcal{A}$, and thus achieves true scalability.
VII-D Sensitivity Analysis and Sampling Frequency Threshold
@@
To better understand the relationship between $h^{\star}$ and the maximum rate constraint parameter $f_{\max}$, we conduct a sensitivity analysis in this subsection to explore how $h^{\star}$ depends on the maximum rate constraint with parameter $f_{\max}$. The key result is stated in the following Lemma, whose proof can be found in Appendix O.
@@
(Sensitivity Analysis). Define $f_{\max}^{T}$ as: $f_{\max}^{T}\triangleq\frac{1}{\mathcal{F}^{(\rho^{\star})^{-}}},$ the following assertions hold true:
(ii). If $f_{\max}\geq f_{\max}^{T}$, $h^{\star}$ is independent of $f_{\max}$;
(iiii). If $f_{\max}<f_{\max}^{T}$, $h^{\star}=f_{\max}Q^{\star}(f_{\max})$ is monotonically non-increasing with $f_{\max}$, with the derivative given as: $\frac{\mathrm{d}h^{\star}}{\mathrm{d}f_{\max}}=U(\lambda^{\star}),$ where $\lambda^{\star}=\underset{\lambda}{\arg\max}\left\{\lambda+f_{\max}U(\lambda)\right\}.$
@@ VIII-A1 Sampling Policies
In this paper, the sampling policy that is co-designed with the goal-oriented remote decision-making policy to minimize the long-term average cost in Problem 1 is referred to as “Goal-oriented sampling”. We compare this sampling policy with the following benchmarks:
  • Uniform sampling: The sampling is activated periodically with a constant sampling interval $d\in\mathbb{N}^{+}$. In this case, the sampling process follows $S_{i+1}-S_{i}=d$ for $\forall i$ (see [43, Section VI] for a detailed system description of uniform sampling with queuing and random service delay). Since the sampling interval $d$ is limited to integer values in the discrete-time MDP setting, the corresponding sampling frequency $1/d$ also takes discrete values. Hence, the simulation yields a performance curve comprising discrete points, as shown in Fig. 10.
@@
  • AoI-optimal sampling: The AoI-optimal policy determines $Z_{i}$ by [88, Theorem 4], which is a threshold-based policy $Z_{i}=\max(0,\beta-Y_{i})$, where $\beta$ is the solution to the following equations:
@@
    with $z(Y)=\max(0,\beta-Y).$ [88, Algorithm 2] presents a low-complexity algorithm to solve $\beta$.
@@ VIII-A2 Remote Decision Making
We consider the following decision-making policies:
@@
  • Long-term optimal decision policy $\pi^{\star}$: The long-term optimal policy thinks ahead to minimize the long-term average cost over time. We compute it by solving the ACOE of the primal MDP using the RVI algorithm. The resulting policy employed at the remote decision maker $\pi^{\star}:\mathcal{S}\rightarrow\mathcal{A}$ selects an action based on the latest available observation $X_{t-\Delta(t)}$, according to $a_{t}=\pi^{\star}(X_{t-\Delta(t)})$. Unless otherwise stated, this long-term optimal decision policy is employed as the default downstream decision-making policy under all benchmark sampling strategies. Throughout the paper, each baseline (e.g., ”zero-wait”, ”AoI-optimal”, ”constant-wait”) refers to a composite strategy that combines the corresponding sampling policy with this long-term optimal decision policy.
Figure 7: Average cost vs. Average delay $\mathbb{E}[Y_{i}]$ under binary delay model ($Y_{\max}=11$), with parameter $p$ controlling the average delay. All baseline sampling policies are paired with the corresponding long-term optimal decision policy.
Figure 8: Average cost vs. Average delay $\mathbb{E}[Y_{i}]$ under binary delay model ($p=0.3$), with $Y_{\max}$ controlling the average delay. All baseline sampling policies are paired with the corresponding long-term optimal decision policy.
Figure 9: Average cost vs. Cutoff $Y_{\max}$ under truncated geometric delay model with $q=0.3$. All baseline sampling policies are paired with the corresponding long-term optimal decision policy.
Figure 10: Average cost vs. Maximum sampling frequency $f_{\max}$. All baseline sampling policies are paired with the corresponding long-term optimal decision policy.
@@ VIII-B1 Primal MDP
Figure 11: Simulation Setup: Transition diagram of the primal MDP given in Appendix H.
As a case study, we specifically consider a benchmark setup for clarity and insight. In this setup, the sizes of the state space and the action space of the primal MDP are both $2$, consisting of states $s_{0},s_{1}$ and actions $a_{0},a_{1}$ [^13] The binary state space is broadly representative of many real-world systems. For instance, it can model the occurrence or absence of critical events such as fires, industrial failures, security breaches, or anomalies in equipment status. The binary action space can be interpreted as whether or not to apply a control intervention in response to the critical event, e.g. extinguishing a detected fire or repairing a malfunctioning piece of equipment.. The primal MDP tuple is detailed in Appendix H, and is visualized by transition diagram given in Fig. 11.
@@
indicating a constant action regardless of state. Consequently, this decision-making policy does not utilize state information at the remote decision maker, and thus the value of information transmission is null in this context.
@@
which clearly depends on the system state. In this case, the remote decision maker must rely on the potentially stale state to select the “right” action, highlighting the utility of information transmission under the decision-making policy.
VIII-B2 Finite-support Memory-less Delays
@@
Throughout, we restrict attention to delay distributions with finite support. It is also practically justified, as many communication and control systems enforce a maximum admissible delay (via timeouts, buffer limits, or QoS deadlines). As a representative example, we consider a binary-valued random delay model where each delay $Y_{i}$ takes value $1$ with probability $p$ and $Y_{\max}$ with probability 1p1-p, i.e., $\Pr(Y_{i}=1)=p,\Pr(Y_{i}=Y_{\max})=1-p,\forall i\in\mathbb{Z}^{+}$ (cf. [9]). This binary-delay formulation enables tractable analysis while capturing key aspects of stochastic delay.
Importantly, the proposed policy maps $(X_{S_{i}},Y_{i},A_{i-1})$ to actions, so the solution complexity scales with $|\mathcal{Y}|$; a finite-support assumption keeps the policy solution computationally manageable. The analysis extends to any finite-support distribution without memory, including truncated variants of heavy-tailed models [^14] Beyond truncation, handling unbounded-support delays (e.g., geometric channels) directly would require additional conditions to ensure average-cost optimality; a systematic treatment is an interesting direction for future work.. Specifically, a geometric channel can be approximated by a truncated geometric delay with parameter $q\in(0,1)$ and cutoff $Y_{\max}$:
$\Pr(Y=y)=\frac{q(1-q)^{y-1}}{1-(1-q)^{Y_{\max}}},\quad y=1,\dots,Y_{\max}.$ (72)
VIII-C Discussions
TABLE III: Relative cost reduction (%) of goal-oriented sampling compared with baselines at selected average delays.
$\mathbb{E}[Y_{i}]$ Zero-wait AoI-optimal Const-wait $z=2$
1.7 4.18% 4.18% 9.98%
5.9 6.23% 6.85% 6.09%
8.0 7.18% 7.83% 6.66%
14.3 10.11% 9.87% 8.76%
Fig. 7 and Fig. 8 illustrate the performance comparisons between various sampling and decision-making policies under different average delays $\mathbb{E}[Y_{i}]$. In Fig. 7, the average delay is adjusted by adjusting $p$, with the maximum delay fixed at $Y_{\max}=11$. In Fig. 8, the average delay is increased by enlarging $Y_{\max}$ while keeping $p=0.3$ fixed. Across both scenarios, the proposed goal-oriented sampling strategy consistently outperforms all baseline methods in terms of minimizing the average cost. Table III reports the relative cost reduction $\eta_{b}=(C_{b}-\rho^{\star})/C_{b}\times 100\%$ of the proposed goal-oriented sampling compared with representative baselines at several average delay levels $\mathbb{E}[Y_{i}]$, where $C_{b}$ is the average cost of baseline $b$. The average delay is controlled by adjusting $Y_{\max}$. For small delays, goal-oriented sampling achieves about 4% reduction over zero-wait and AoI-optimal policies and nearly 10% over the constant-wait policy. As the delay increases to moderate levels ($\mathbb{E}[Y_{i}]=5.9$88), the improvement grows to 6–8%, and at large delays ($\mathbb{E}[Y_{i}]=14.3$) the gain further increases to about 10% against zero-wait and AoI-optimal. The performance gain primarily arises from the fact that the proposed sampling policy is explicitly designed to optimize decision effectiveness rather than communication metrics alone. Unlike conventional AoI-optimal sampling policy, which treats information freshness and control decisions as separate optimization problems, goal-oriented sampling allocates sampling opportunities more judiciously, prioritizing updates that are most relevant to improving decision outcomes. In contrast, AoI-optimal sampling may transmit timely but less informative data, resulting in suboptimal decision performance despite lower AoI.